# Observable Primordial Gravitational Waves from Non-minimally Coupled $R^{2}$ Palatini Modified Gravity

### v1

@@ Appendix C Solving for Mode Functions
$v(x)=\sqrt{x}\left(c_{1}H_{\nu}^{(1)}(x)+c_{2}H_{\nu}^{(2)}(x)\right)$ (C.1)
@@
$H_{\nu}^{(1)}(x)\sim\sqrt{\frac{2}{\pi x}}e^{i\left(x-\frac{\nu\pi}{2}-\frac{\pi}{4}\right)}\hskip 8.5359pt\text{and}\hskip 8.5359ptH_{\nu}^{(2)}(x)\sim\sqrt{\frac{2}{\pi x}}e^{-i\left(x-\frac{\nu\pi}{2}-\frac{\pi}{4}\right)}$ (C.2)
@@
$\lim_{\eta\to-\infty}v(\eta,k)=\frac{1}{\sqrt{2k}}e^{-ik\eta}$ (C.3)
@@
$v(x)=c_{1}\sqrt{x}H_{\nu_{I}}^{(1)}(x)$ (C.4)
In the sub-Hubble regime ($x\to\infty$), the Hankel function of the first kind takes the form of the first equation in (C.2). Thus for $x\gg 1$, the mode function becomes,
$v(x)\sim c_{1}\sqrt{\frac{2}{\pi}}e^{i\left(x-\frac{\nu_{I}\pi}{2}-\frac{\pi}{4}\right)}$ (C.5)
@@
$c_{1}\sqrt{\frac{2}{\pi}}=\frac{1}{\sqrt{2k}}e^{\frac{i\pi}{4}(1+2\nu_{I})}\implies c_{1}=\sqrt{\frac{\pi}{4k}}e^{\frac{i\pi}{4}(1+2\nu_{I})}$ (C.7)
@@
$\begin{split}v(x)&=\sqrt{\frac{\pi}{4k}}\sqrt{x}~{}e^{\frac{i\pi}{4}(1+2\nu_{I})}H_{\nu_{I}}^{(1)}(x)\end{split}$ (C.8)
$\begin{split}v_{k}^{s}(\eta)&=\sqrt{\frac{\pi}{4k}}\sqrt{-k\eta}~{}e^{\frac{i\pi}{4}(1+2\nu_{I})}H_{\nu_{I}}^{(1)}(-k\eta)\end{split}$ (C.9)
@@
$v_{k}^{s}(\eta)=\frac{1}{\sqrt{2k}}\left[\alpha_{+}(k)e^{-ik\eta}+\alpha_{-}(k)e^{ik\eta}\right]$ (C.10)
A similar solution with different coefficients ($\gamma_{\pm}$) follows for the era of radiation-domination. We now focus on the era of kination. This phase does not require a Bunch-Davies vacuum condition and thus can have negative-frequency modes. Therefore the solution is similar to (C.9),
$\begin{split}v_{k}^{s}(\eta)&=\sqrt{\frac{\pi}{4k}}\sqrt{kz}~{}e^{\frac{i\pi}{4}(1+2\nu_{II})}H_{\nu_{II}}^{(1)}(kz)+\sqrt{\frac{\pi}{4k}}\sqrt{kz}~{}e^{-\frac{i\pi}{4}(1+2\nu_{II})}H_{\nu_{II}}^{(2)}(kz)\end{split}$ (C.11)
$\begin{split}&=\sqrt{\frac{\pi z}{4}}\left[\beta_{+}(k)~{}e^{-\frac{i\pi}{4}(1+2\nu_{II})}H_{\nu_{II}}^{(2)}(kz)+\beta_{-}(k)~{}e^{\frac{i\pi}{4}(1+2\nu_{II})}H_{\nu_{II}}^{(1)}(kz)\right]\end{split}$ (C.12)
@@
Appendix D Matching the Mode Functions
@@
$v^{s^{\prime\prime}}_{k}+k^{2}v^{s}_{k}=0\implies v^{s}_{k}(\eta)=\frac{1}{\sqrt{2k}}\left[\alpha_{+}(k)e^{-ik\eta}+\alpha_{-}(k)e^{ik\eta}\right]$ (D.1)
@@
$\sqrt{\frac{\pi}{2}}\sqrt{x_{\text{end}}}~{}e^{\frac{i\pi}{4}(1+2\nu_{I})}H_{\nu_{I}}^{(1)}(x_{\text{end}})=\alpha_{+}e^{ik|\eta_{\text{end}}|}+\alpha_{-}e^{-ik|\eta_{\text{end}}|}$ (D.2)
@@
$ie^{\frac{i\pi}{4}(1+2\nu_{I})}\sqrt{\frac{\pi}{2}}\left[\frac{1}{\sqrt{x_{\text{end}}}}\left(\frac{1}{2}+\nu\right)H_{\nu_{I}}^{(1)}(x_{\text{end}})-\sqrt{x_{\text{end}}}H_{\nu_{I}+1}^{(1)}(x_{\text{end}})\right]=-\alpha_{+}e^{ik|\eta_{\text{end}}|}+\alpha_{-}e^{-ik|\eta_{\text{end}}|}$ (D.3)
Adding and subtracting (D.2) and (D.3) gives us,
$\alpha_{\mp}=\frac{1}{2}e^{\frac{i\pi}{4}(1+2\nu_{I})\pm ix_{\text{end}}}\sqrt{\frac{\pi}{2}}\left[H_{\nu_{I}}^{(1)}(x_{\text{end}})\left(\sqrt{x_{\text{end}}}\pm i\frac{1}{\sqrt{x_{\text{end}}}}\left(\frac{1}{2}+\nu\right)\right)\mp i\sqrt{x_{\text{end}}}H_{\nu_{I}+1}^{(1)}(x_{\text{end}})\right]$ (D.4)
@@
$\alpha_{\mp}$ $=\frac{1}{2}e^{\frac{i\pi}{4}(1+2\nu_{I})}\sqrt{\frac{\pi}{2}}\left[H_{\nu_{I}}^{(1)}(x_{\text{end}})\left(\pm\frac{i}{\sqrt{x_{\text{end}}}}\left(\frac{1}{2}+\nu\right)\right)\mp i\sqrt{x_{\text{end}}}H_{\nu_{I}+1}^{(1)}(x_{\text{end}})\right]$ (D.5)
@@
$\alpha_{\mp}=\pm 2^{\nu_{I}-1}e^{\frac{i\pi}{4}(1+2\nu_{I})}\sqrt{\frac{1}{2\pi}}\left(\frac{1}{2}-\nu\right)\Gamma(\nu_{I})\frac{1}{x_{\text{end}}^{\nu_{I}+1/2}}$ (D.6)
@@
$\alpha_{\mp}$ =±$=\pm 2^{\epsilon-1}e^{\frac{i\pi}{2}\epsilon}(\epsilon+1)\frac{\Gamma(3/2+\epsilon)}{\Gamma(3/2)}\left(\frac{H}{k}\right)^{2+\epsilon}$ (D.7)
@@
$\alpha_{\mp}=\pm\frac{H^{2}}{2k^{2}}$ (D.8)
The next transition occurs at $\eta_{\text{kin}}$, for the phase moving from hyperkination to kination. Solving the Mukhanov-Sasaki equation as done in Appendix C, we use (C.11) with the redefinition of $y\equiv kz$ and $\nu_{II}=1/2-1/p$.
$v_{k}^{s}(\eta)=\sqrt{\frac{\pi}{4k}}\sqrt{y}\left[\beta_{-}(k)~{}e^{\frac{i\pi}{4}(1+2\nu_{II})}H_{\nu_{II}}^{(1)}(y)+\beta_{+}(k)~{}e^{-\frac{i\pi}{4}(1+2\nu_{II})}H_{\nu_{II}}^{(2)}(y)\right]$ (D.9)
@@
$y_{\text{kin}}\equiv kz_{\text{kin}}=k\left(\frac{\eta_{\text{kin}}}{2}+\frac{1}{H}\right)\approx\frac{k\eta_{\text{kin}}}{2}$ (D.10)
@@
$r\equiv\sqrt{\frac{\pi}{2}}~{}e^{\frac{i\pi}{4}(1+2\nu_{II})}\quad\text{and}\quad r^{\ast}\equiv\sqrt{\frac{\pi}{2}}~{}e^{\frac{-i\pi}{4}(1+2\nu_{II})}$ (D.11)
@@
$\alpha_{+}e^{-ik\eta_{\text{kin}}}+\alpha_{-}e^{ik\eta_{\text{kin}}}=\sqrt{y_{\text{kin}}}\left[\beta_{-}rH_{\nu_{II}}^{(1)}(y_{\text{kin}})+\beta_{+}r^{\ast}H_{\nu_{II}}^{(2)}(y_{\text{kin}})\right]$ (D.12)
@@
$\begin{split}i\left(-\alpha_{+}e^{-ik\eta_{\text{kin}}}+\alpha_{-}e^{ik\eta_{\text{kin}}}\right)&=\frac{1}{2\sqrt{y_{\text{kin}}}}\left[\beta_{-}rH_{\nu_{II}}^{(1)}(y_{\text{kin}})+\beta_{+}r^{\ast}H_{\nu_{II}}^{(2)}(y_{\text{kin}})\right]\\ &+\sqrt{y_{\text{kin}}}\left[\beta_{-}r\frac{\mathrm{d}H_{\nu_{II}}^{(1)}}{\mathrm{d}y}(y_{\text{kin}})+\beta_{+}r^{\ast}\frac{\mathrm{d}H_{\nu_{II}}^{(2)}}{\mathrm{d}y}(y_{\text{kin}})\right]\end{split}$ (D.13)
Using (D.12),
$\begin{split}\alpha_{+}\left(-i-\frac{1}{2y_{\text{kin}}}\right)e^{-ik\eta_{\text{kin}}}+\alpha_{-}\left(i-\frac{1}{2y_{\text{kin}}}\right)e^{ik\eta_{\text{kin}}}&=\sqrt{y_{\text{kin}}}\left[\beta_{-}r\frac{\mathrm{d}H_{\nu_{II}}^{(1)}}{\mathrm{d}y}(y_{\text{kin}})+\beta_{+}r^{\ast}\frac{\mathrm{d}H_{\nu_{II}}^{(2)}}{\mathrm{d}y}(y_{\text{kin}})\right]\end{split}$ (D.14)
To get $\beta_{-}$, we multiply (D.14) by $H_{\nu_{II}}^{(2)}(y_{\text{kin}})$ and (D.12) by $\mathrm{d}H_{\nu_{II}}^{(2)}/\mathrm{d}y$ and subtract the latter from the former and use the Wronskian of the Hankel functions.
$\begin{split}H_{\nu_{II}}^{(2)}\left(\alpha_{+}\left(-i-\frac{1}{2y_{\text{kin}}}\right)e^{-ik\eta_{\text{kin}}}+\alpha_{-}\left(i-\frac{1}{2y_{\text{kin}}}\right)e^{ik\eta_{\text{kin}}}\right)&=\sqrt{y_{\text{kin}}}\left[\beta_{-}rH_{\nu_{II}}^{(2)}\frac{\mathrm{d}H_{\nu_{II}}^{(1)}}{\mathrm{d}y}(y_{\text{kin}})\right.\\ &\left.+\beta_{+}r^{\ast}H_{\nu_{II}}^{(2)}\frac{\mathrm{d}H_{\nu_{II}}^{(2)}}{\mathrm{d}y}(y_{\text{kin}})\right]\end{split}$ (D.15)
$\begin{split}\beta_{-}&=e^{-\frac{i\pi}{4}(1+2\nu_{II})}\frac{\sqrt{\pi y_{\text{kin}}}}{i2\sqrt{2}}\left[H_{\nu_{II}}^{(2)}\left(\alpha_{+}\left[-i-\frac{1}{2y_{\text{kin}}}\right]e^{-ik\eta_{\text{kin}}}+\alpha_{-}\left[i-\frac{1}{2y_{\text{kin}}}\right]e^{ik\eta_{\text{kin}}}\right)\right.\\ &\left.-\frac{1}{2}\left(H_{\nu_{II}-1}^{(2)}-H_{\nu_{II}+1}^{(2)}\right)\left(\alpha_{+}e^{-ik\eta_{\text{kin}}}+\alpha_{-}e^{ik\eta_{\text{kin}}}\right)\right]\end{split}$ (D.16)
Now to get $\beta_{+}$, we multiply (D.14) by $H_{\nu_{II}}^{(1)}(y_{\text{kin}})$ and (D.12) by $\mathrm{d}H_{\nu_{II}}^{(1)}/\mathrm{d}y$ and follow the same procedure, we get,
$\begin{split}\beta_{+}&=-e^{\frac{i\pi}{4}(1+2\nu_{II})}\frac{\sqrt{\pi y_{\text{kin}}}}{i2\sqrt{2}}\left[H_{\nu_{II}}^{(1)}\left(\alpha_{+}\left[-i-\frac{1}{2y_{\text{kin}}}\right]e^{-ik\eta_{\text{kin}}}+\alpha_{-}\left[i-\frac{1}{2y_{\text{kin}}}\right]e^{ik\eta_{\text{kin}}}\right)\right.\\ &\left.-\frac{1}{2}\left(H_{\nu_{II}-1}^{(1)}-H_{\nu_{II}+1}^{(1)}\right)\left(\alpha_{+}e^{-ik\eta_{\text{kin}}}+\alpha_{-}e^{ik\eta_{\text{kin}}}\right)\right]\end{split}$ (D.17)
@@
$\begin{split}\beta_{-}&=e^{-\frac{i\pi}{4}(1+2\nu_{II})}\frac{\sqrt{\pi y_{\text{kin}}}}{i2\sqrt{2}}\left[H_{\nu_{II}}^{(2)}\left(i\alpha_{-}\left(e^{-ik\eta_{\text{kin}}}+e^{ik\eta_{\text{kin}}}\right)+\frac{\alpha_{-}}{2y_{\text{kin}}}\left(e^{-ik\eta_{\text{kin}}}-e^{ik\eta_{\text{kin}}}\right)\right)\right.\\ &\left.-\frac{\alpha_{-}}{2}\left(H_{\nu_{II}-1}^{(2)}-H_{\nu_{II}+1}^{(2)}\right)\left(-e^{-ik\eta_{\text{kin}}}+e^{ik\eta_{\text{kin}}}\right)\right]\end{split}$ (D.18)
$\begin{split}\beta_{-}&=e^{-\frac{i\pi}{4}(1+2\nu_{II})}\sqrt{\frac{\pi y_{\text{kin}}}{2}}~{}\alpha_{-}\left[H_{\nu_{II}}^{(2)}\left(\cos{(k\eta_{\text{kin}})}-\frac{1}{2y_{\text{kin}}}\sin{(k\eta_{\text{kin}})}\right)\right.\\ &\left.-\left(H_{\nu_{II}-1}^{(2)}-H_{\nu_{II}+1}^{(2)}\right)\sin{(k\eta_{\text{kin}})}\right]\end{split}$ (D.19)
@@
$\begin{split}\beta_{+}&=-e^{\frac{i\pi}{4}(1+2\nu_{II})}\sqrt{\frac{\pi y_{\text{kin}}}{2}}~{}\alpha_{-}\left[H_{\nu_{II}}^{(1)}\left(\cos{(k\eta_{\text{kin}})}-\frac{1}{2y_{\text{kin}}}\sin{(k\eta_{\text{kin}})}\right)\right.\\ &\left.-\left(H_{\nu_{II}-1}^{(1)}-H_{\nu_{II}+1}^{(1)}\right)\sin{(k\eta_{\text{kin}})}\right]\end{split}$ (D.20)
@@
$\beta_{-}$ $=e^{-\frac{i\pi}{4}(1+2\nu_{II})}\sqrt{\frac{\pi y_{\text{kin}}}{2}}~{}\alpha_{-}\left[-\left(H_{\nu_{II}-1}^{(2)}-H_{\nu_{II}+1}^{(2)}\right)k\eta_{\text{kin}}\right]$ (D.21)
$\beta_{-}$ $=-e^{-\frac{i\pi}{4}(1+2\nu_{II})}\sqrt{\frac{\pi}{4}}~{}\frac{H^{2}}{2k^{2}}\cdot 2\left.\frac{\mathrm{d}H_{\nu_{II}}^{(2)}}{\mathrm{d}y}\right|_{y_{\text{kin}}}(k\eta_{\text{kin}})^{3/2}$ (D.22)
$\begin{split}&=e^{-\frac{i\pi}{4}(1+2\nu_{II})}\sqrt{\frac{\pi}{4}}~{}\frac{H^{2}}{2k^{2}}\cdot 2\frac{i}{\pi}\Gamma(\nu_{II}+1)\left(\frac{y_{\text{kin}}}{2}\right)^{-\nu_{II}-1}(k\eta_{\text{kin}})^{3/2}\end{split}$ (D.23)
$\beta_{-}=ie^{-\frac{i\pi}{4}(1+2\nu_{II})}2^{2\nu_{II}}~{}\frac{H^{2}}{k^{2}}\frac{\Gamma(3/2-1/p)}{\Gamma(3/2)}(k\eta_{\text{kin}})^{1/p}$ (D.24)

### v2

@@ Appendix D Solving for Mode Functions
$v(x)=\sqrt{x}\left(c_{1}H_{\nu}^{(1)}(x)+c_{2}H_{\nu}^{(2)}(x)\right)$ (D.1)
@@
$H_{\nu}^{(1)}(x)\sim\sqrt{\frac{2}{\pi x}}e^{i\left(x-\frac{\nu\pi}{2}-\frac{\pi}{4}\right)}\hskip 8.5359pt\text{and}\hskip 8.5359ptH_{\nu}^{(2)}(x)\sim\sqrt{\frac{2}{\pi x}}e^{-i\left(x-\frac{\nu\pi}{2}-\frac{\pi}{4}\right)}$ (D.2)
@@
$\lim_{\eta\to-\infty}v(\eta,k)=\frac{1}{\sqrt{2k}}e^{-ik\eta}$ (D.3)
@@
$v(x)=c_{1}\sqrt{x}H_{\nu_{I}}^{(1)}(x)$ (D.4)
In the sub-Hubble regime ($x\to\infty$), the Hankel function of the first kind takes the form of the first equation in (D.2). Thus for $x\gg 1$, the mode function becomes,
$v(x)\sim c_{1}\sqrt{\frac{2}{\pi}}e^{i\left(x-\frac{\nu_{I}\pi}{2}-\frac{\pi}{4}\right)}$ (D.5)
@@
$c_{1}\sqrt{\frac{2}{\pi}}=\frac{1}{\sqrt{2k}}e^{\frac{i\pi}{4}(1+2\nu_{I})}\implies c_{1}=\sqrt{\frac{\pi}{4k}}e^{\frac{i\pi}{4}(1+2\nu_{I})}$ (D.7)
@@
$\begin{split}v(x)&=\sqrt{\frac{\pi}{4k}}\sqrt{x}~{}e^{\frac{i\pi}{4}(1+2\nu_{I})}H_{\nu_{I}}^{(1)}(x)\end{split}$ (D.8)
$\begin{split}v_{k}^{s}(\eta)&=\sqrt{\frac{\pi}{4k}}\sqrt{-k\eta}~{}e^{\frac{i\pi}{4}(1+2\nu_{I})}H_{\nu_{I}}^{(1)}(-k\eta)\end{split}$ (D.9)
@@
$v_{k}^{s}(\eta)=\frac{1}{\sqrt{2k}}\left[\alpha_{+}(k)e^{-ik\eta}+\alpha_{-}(k)e^{ik\eta}\right]$ (D.10)
A similar solution with different coefficients ($\gamma_{\pm}$) follows for the era of radiation-domination. We now focus on the era of kination. This phase does not require a Bunch-Davies vacuum condition and thus can have negative-frequency modes. Therefore the solution is similar to (D.9),
$\begin{split}v_{k}^{s}(\eta)&=\sqrt{\frac{\pi}{4k}}\sqrt{kz}~{}e^{\frac{i\pi}{4}(1+2\nu_{II})}H_{\nu_{II}}^{(1)}(kz)+\sqrt{\frac{\pi}{4k}}\sqrt{kz}~{}e^{-\frac{i\pi}{4}(1+2\nu_{II})}H_{\nu_{II}}^{(2)}(kz)\end{split}$ (D.11)
$\begin{split}&=\sqrt{\frac{\pi z}{4}}\left[\beta_{+}(k)~{}e^{-\frac{i\pi}{4}(1+2\nu_{II})}H_{\nu_{II}}^{(2)}(kz)+\beta_{-}(k)~{}e^{\frac{i\pi}{4}(1+2\nu_{II})}H_{\nu_{II}}^{(1)}(kz)\right]\end{split}$ (D.12)
@@
Appendix E Matching the Mode Functions
@@
$v^{s^{\prime\prime}}_{k}+k^{2}v^{s}_{k}=0\implies v^{s}_{k}(\eta)=\frac{1}{\sqrt{2k}}\left[\alpha_{+}(k)e^{-ik\eta}+\alpha_{-}(k)e^{ik\eta}\right]$ (E.1)
@@
$\sqrt{\frac{\pi}{2}}\sqrt{x_{\text{end}}}~{}e^{\frac{i\pi}{4}(1+2\nu_{I})}H_{\nu_{I}}^{(1)}(x_{\text{end}})=\alpha_{+}e^{ik|\eta_{\text{end}}|}+\alpha_{-}e^{-ik|\eta_{\text{end}}|}$ (E.2)
@@
$ie^{\frac{i\pi}{4}(1+2\nu_{I})}\sqrt{\frac{\pi}{2}}\left[\frac{1}{\sqrt{x_{\text{end}}}}\left(\frac{1}{2}+\nu\right)H_{\nu_{I}}^{(1)}(x_{\text{end}})-\sqrt{x_{\text{end}}}H_{\nu_{I}+1}^{(1)}(x_{\text{end}})\right]=-\alpha_{+}e^{ik|\eta_{\text{end}}|}+\alpha_{-}e^{-ik|\eta_{\text{end}}|}$ (E.3)
Adding and subtracting (E.2) and (E.3) gives us,
$\alpha_{\mp}=\frac{1}{2}e^{\frac{i\pi}{4}(1+2\nu_{I})\pm ix_{\text{end}}}\sqrt{\frac{\pi}{2}}\left[H_{\nu_{I}}^{(1)}(x_{\text{end}})\left(\sqrt{x_{\text{end}}}\pm i\frac{1}{\sqrt{x_{\text{end}}}}\left(\frac{1}{2}+\nu\right)\right)\mp i\sqrt{x_{\text{end}}}H_{\nu_{I}+1}^{(1)}(x_{\text{end}})\right]$ (E.4)
@@
$\alpha_{\mp}$ $=\frac{1}{2}e^{\frac{i\pi}{4}(1+2\nu_{I})}\sqrt{\frac{\pi}{2}}\left[H_{\nu_{I}}^{(1)}(x_{\text{end}})\left(\pm\frac{i}{\sqrt{x_{\text{end}}}}\left(\frac{1}{2}+\nu\right)\right)\mp i\sqrt{x_{\text{end}}}H_{\nu_{I}+1}^{(1)}(x_{\text{end}})\right]$ (E.5)
@@
$\alpha_{\mp}=\pm 2^{\nu_{I}-1}e^{\frac{i\pi}{4}(1+2\nu_{I})}\sqrt{\frac{1}{2\pi}}\left(\frac{1}{2}-\nu\right)\Gamma(\nu_{I})\frac{1}{x_{\text{end}}^{\nu_{I}+1/2}}$ (E.6)
@@
$\alpha_{\mp}$ =±$=\pm 2^{\epsilon-1}e^{\frac{i\pi}{2}\epsilon}(\epsilon+1)\frac{\Gamma(3/2+\epsilon)}{\Gamma(3/2)}\left(\frac{H}{k}\right)^{2+\epsilon}$ (E.7)
@@
$\alpha_{\mp}=\pm\frac{H^{2}}{2k^{2}}$ (E.8)
The next transition occurs at $\eta_{\text{kin}}$, for the phase moving from hyperkination to kination. Solving the Mukhanov-Sasaki equation as done in Appendix D, we use (D.11) with the redefinition of $y\equiv kz$ and $\nu_{II}=1/2-1/p$.
$v_{k}^{s}(\eta)=\sqrt{\frac{\pi}{4k}}\sqrt{y}\left[\beta_{-}(k)~{}e^{\frac{i\pi}{4}(1+2\nu_{II})}H_{\nu_{II}}^{(1)}(y)+\beta_{+}(k)~{}e^{-\frac{i\pi}{4}(1+2\nu_{II})}H_{\nu_{II}}^{(2)}(y)\right]$ (E.9)
@@
$y_{\text{kin}}\equiv kz_{\text{kin}}=k\left(\frac{\eta_{\text{kin}}}{2}+\frac{1}{H}\right)\approx\frac{k\eta_{\text{kin}}}{2}$ (E.10)
@@
$r\equiv\sqrt{\frac{\pi}{2}}~{}e^{\frac{i\pi}{4}(1+2\nu_{II})}\quad\text{and}\quad r^{\ast}\equiv\sqrt{\frac{\pi}{2}}~{}e^{\frac{-i\pi}{4}(1+2\nu_{II})}$ (E.11)
@@
$\alpha_{+}e^{-ik\eta_{\text{kin}}}+\alpha_{-}e^{ik\eta_{\text{kin}}}=\sqrt{y_{\text{kin}}}\left[\beta_{-}rH_{\nu_{II}}^{(1)}(y_{\text{kin}})+\beta_{+}r^{\ast}H_{\nu_{II}}^{(2)}(y_{\text{kin}})\right]$ (E.12)
@@
$\begin{split}i\left(-\alpha_{+}e^{-ik\eta_{\text{kin}}}+\alpha_{-}e^{ik\eta_{\text{kin}}}\right)&=\frac{1}{2\sqrt{y_{\text{kin}}}}\left[\beta_{-}rH_{\nu_{II}}^{(1)}(y_{\text{kin}})+\beta_{+}r^{\ast}H_{\nu_{II}}^{(2)}(y_{\text{kin}})\right]\\ &+\sqrt{y_{\text{kin}}}\left[\beta_{-}r\frac{\mathrm{d}H_{\nu_{II}}^{(1)}}{\mathrm{d}y}(y_{\text{kin}})+\beta_{+}r^{\ast}\frac{\mathrm{d}H_{\nu_{II}}^{(2)}}{\mathrm{d}y}(y_{\text{kin}})\right]\end{split}$ (E.13)
Using (E.12),
$\begin{split}\alpha_{+}\left(-i-\frac{1}{2y_{\text{kin}}}\right)e^{-ik\eta_{\text{kin}}}+\alpha_{-}\left(i-\frac{1}{2y_{\text{kin}}}\right)e^{ik\eta_{\text{kin}}}&=\sqrt{y_{\text{kin}}}\left[\beta_{-}r\frac{\mathrm{d}H_{\nu_{II}}^{(1)}}{\mathrm{d}y}(y_{\text{kin}})+\beta_{+}r^{\ast}\frac{\mathrm{d}H_{\nu_{II}}^{(2)}}{\mathrm{d}y}(y_{\text{kin}})\right]\end{split}$ (E.14)
To get $\beta_{-}$, we multiply (E.14) by $H_{\nu_{II}}^{(2)}(y_{\text{kin}})$ and (E.12) by $\mathrm{d}H_{\nu_{II}}^{(2)}/\mathrm{d}y$ and subtract the latter from the former and use the Wronskian of the Hankel functions.
$\begin{split}H_{\nu_{II}}^{(2)}\left(\alpha_{+}\left(-i-\frac{1}{2y_{\text{kin}}}\right)e^{-ik\eta_{\text{kin}}}+\alpha_{-}\left(i-\frac{1}{2y_{\text{kin}}}\right)e^{ik\eta_{\text{kin}}}\right)&=\sqrt{y_{\text{kin}}}\left[\beta_{-}rH_{\nu_{II}}^{(2)}\frac{\mathrm{d}H_{\nu_{II}}^{(1)}}{\mathrm{d}y}(y_{\text{kin}})\right.\\ &\left.+\beta_{+}r^{\ast}H_{\nu_{II}}^{(2)}\frac{\mathrm{d}H_{\nu_{II}}^{(2)}}{\mathrm{d}y}(y_{\text{kin}})\right]\end{split}$ (E.15)
$\begin{split}\beta_{-}&=e^{-\frac{i\pi}{4}(1+2\nu_{II})}\frac{\sqrt{\pi y_{\text{kin}}}}{i2\sqrt{2}}\left[H_{\nu_{II}}^{(2)}\left(\alpha_{+}\left[-i-\frac{1}{2y_{\text{kin}}}\right]e^{-ik\eta_{\text{kin}}}+\alpha_{-}\left[i-\frac{1}{2y_{\text{kin}}}\right]e^{ik\eta_{\text{kin}}}\right)\right.\\ &\left.-\frac{1}{2}\left(H_{\nu_{II}-1}^{(2)}-H_{\nu_{II}+1}^{(2)}\right)\left(\alpha_{+}e^{-ik\eta_{\text{kin}}}+\alpha_{-}e^{ik\eta_{\text{kin}}}\right)\right]\end{split}$ (E.16)
Now to get $\beta_{+}$, we multiply (E.14) by $H_{\nu_{II}}^{(1)}(y_{\text{kin}})$ and (E.12) by $\mathrm{d}H_{\nu_{II}}^{(1)}/\mathrm{d}y$ and follow the same procedure, we get,
$\begin{split}\beta_{+}&=-e^{\frac{i\pi}{4}(1+2\nu_{II})}\frac{\sqrt{\pi y_{\text{kin}}}}{i2\sqrt{2}}\left[H_{\nu_{II}}^{(1)}\left(\alpha_{+}\left[-i-\frac{1}{2y_{\text{kin}}}\right]e^{-ik\eta_{\text{kin}}}+\alpha_{-}\left[i-\frac{1}{2y_{\text{kin}}}\right]e^{ik\eta_{\text{kin}}}\right)\right.\\ &\left.-\frac{1}{2}\left(H_{\nu_{II}-1}^{(1)}-H_{\nu_{II}+1}^{(1)}\right)\left(\alpha_{+}e^{-ik\eta_{\text{kin}}}+\alpha_{-}e^{ik\eta_{\text{kin}}}\right)\right]\end{split}$ (E.17)
@@
$\begin{split}\beta_{-}&=e^{-\frac{i\pi}{4}(1+2\nu_{II})}\frac{\sqrt{\pi y_{\text{kin}}}}{i2\sqrt{2}}\left[H_{\nu_{II}}^{(2)}\left(i\alpha_{-}\left(e^{-ik\eta_{\text{kin}}}+e^{ik\eta_{\text{kin}}}\right)+\frac{\alpha_{-}}{2y_{\text{kin}}}\left(e^{-ik\eta_{\text{kin}}}-e^{ik\eta_{\text{kin}}}\right)\right)\right.\\ &\left.-\frac{\alpha_{-}}{2}\left(H_{\nu_{II}-1}^{(2)}-H_{\nu_{II}+1}^{(2)}\right)\left(-e^{-ik\eta_{\text{kin}}}+e^{ik\eta_{\text{kin}}}\right)\right]\end{split}$ (E.18)
$\begin{split}\beta_{-}&=e^{-\frac{i\pi}{4}(1+2\nu_{II})}\sqrt{\frac{\pi y_{\text{kin}}}{2}}~{}\alpha_{-}\left[H_{\nu_{II}}^{(2)}\left(\cos{(k\eta_{\text{kin}})}-\frac{1}{2y_{\text{kin}}}\sin{(k\eta_{\text{kin}})}\right)\right.\\ &\left.-\left(H_{\nu_{II}-1}^{(2)}-H_{\nu_{II}+1}^{(2)}\right)\sin{(k\eta_{\text{kin}})}\right]\end{split}$ (E.19)
@@
$\begin{split}\beta_{+}&=-e^{\frac{i\pi}{4}(1+2\nu_{II})}\sqrt{\frac{\pi y_{\text{kin}}}{2}}~{}\alpha_{-}\left[H_{\nu_{II}}^{(1)}\left(\cos{(k\eta_{\text{kin}})}-\frac{1}{2y_{\text{kin}}}\sin{(k\eta_{\text{kin}})}\right)\right.\\ &\left.-\left(H_{\nu_{II}-1}^{(1)}-H_{\nu_{II}+1}^{(1)}\right)\sin{(k\eta_{\text{kin}})}\right]\end{split}$ (E.20)
@@
$\beta_{-}$ $=e^{-\frac{i\pi}{4}(1+2\nu_{II})}\sqrt{\frac{\pi y_{\text{kin}}}{2}}~{}\alpha_{-}\left[-\left(H_{\nu_{II}-1}^{(2)}-H_{\nu_{II}+1}^{(2)}\right)k\eta_{\text{kin}}\right]$ (E.21)
$\beta_{-}$ $=-e^{-\frac{i\pi}{4}(1+2\nu_{II})}\sqrt{\frac{\pi}{4}}~{}\frac{H^{2}}{2k^{2}}\cdot 2\left.\frac{\mathrm{d}H_{\nu_{II}}^{(2)}}{\mathrm{d}y}\right|_{y_{\text{kin}}}(k\eta_{\text{kin}})^{3/2}$ (E.22)
$\begin{split}&=e^{-\frac{i\pi}{4}(1+2\nu_{II})}\sqrt{\frac{\pi}{4}}~{}\frac{H^{2}}{2k^{2}}\cdot 2\frac{i}{\pi}\Gamma(\nu_{II}+1)\left(\frac{y_{\text{kin}}}{2}\right)^{-\nu_{II}-1}(k\eta_{\text{kin}})^{3/2}\end{split}$ (E.23)
$\beta_{-}=ie^{-\frac{i\pi}{4}(1+2\nu_{II})}2^{2\nu_{II}}~{}\frac{H^{2}}{k^{2}}\frac{\Gamma(3/2-1/p)}{\Gamma(3/2)}(k\eta_{\text{kin}})^{1/p}$ (E.24)